# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

### v2

@@ 1. Introduction
Ultrametricity in physics means the emergence of ultrametric spaces in physical models. A metric space $(M,d)$ is ultrametric if the distance $d$ satisfies $d(A,B)\leq\max\left\{d(A,C),d(C,B)\right\}$ for any three points $A$, $B,$ $C$ in $M$. Ultrametricity was discovered in the 80s by Parisi and others in the theory of spin glasses and by Frauenfelder and others in physics of proteins; see e.g. [11], [12], see also [13, Chapter 4], and the references therein. In both cases, the space of states of a complex system has a hierarchical structure that plays a central role in the physical behavior of the system.
@@
On the other hand, in the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement satisfies $\Delta x\geq L_{\text{Planck}}:=\sqrt{\frac{\hbar G}{c^{3}}}$, where $L_{\text{Planck}}\approx 10^{-33}$ cmcm is the Planck length. This implies that space-time is not an infinitely divisible continuum, [14]. Mathematically speaking, space-time must be a completely disconnected topological space. The ultrametric spaces are naturally completely disconnected. The field of $p$-adic numbers $\mathbb{Q}_{p}$ is a paramount example of an ultrametric space, with a very rich mathematical structure. There are several possible interpretations of the Bronstein inequality. One of them drives the loop quantum gravity. Volovich gave another interpretation of Bronstein’s inequality in the 80s. The inequality mentioned implies that real numbers cannot be used in models at the level of Planck’s length, because the Archimedean axiom, which appears naturally if we use real numbers, implies that lengths can be measured with arbitrary precision. Volovich proposed using $p$-adic numbers in physical models at the Planck scale, [15], see also [17], [16, Chapter 6].
@@
In the Dirac-von Neumann formulation of quantum mechanics (QM), the states of a closed quantum system are vectors of an abstract complex Hilbert space $\mathcal{H}$, and the observables correspond to linear self-adjoint operators in $\mathcal{H}$. In $p$-adic QM, $\mathcal{H}=L^{2}(\mathbb{Q}_{p}^{N})$ or $\mathcal{H}=L^{2}(U)$, where $U\subset\mathbb{Q}_{p}^{N}$. This choice implies that the position vectors are $\mathbb{Q}_{p}^{N}$, while the time is a real variable. Using $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ as a model of space-time, implies assuming that the space ($\mathbb{Q}_{p}^{3}$) has a discrete nature: there are no continuous word lines joining two different points in the space. The assumption of the discreteness of space, which requires passing from $\mathbb{R}\times\mathbb{R}^{3}$ to $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ (or $\mathbb{R}\times\mathcal{X}$), is incompatible with special and general relativity. We warn the reader that there are several different types of $p$-adic QM; for instance, if the time is assumed to be a $p$-adic variable, the QM obtained radically differs from the one considered here. To the best of our knowledge, $p$-adic QM started in the 1980s under the influence of Vladimirov and Volovich, [18]. The literature about $p$-adic QM is pervasive, and here we cited just a few works, [18]-[26].
@@
In [26], the first author showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, we construct new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. The final goal of the mentioned work was to contribute in the understanding of the role of the hypothesis of the discreteness of space in the foundations of quantum mechanics. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. But, this theory is not a mathematical toy. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
@@
In this work, we study a large class of $p$-adic Schrödinger equations and the associated CTQMCs by comparing them with $p$-adic heat equations and the associated CTMCs. This work is motivated by [5], where a particular class CTQWs was compared with their classical counterparts. Here, the comparison is done by a mathematical study of the mentioned equations, which requires, for instance, solving the initial value problems attached to the mentioned equations, and through numerical simulations. A $p$-adic heat equation is a continuous (or scaling) limit of a CTMC. This assertion has been verified for some classes of equations, see [26], which include the one considered here. But, for general $p$-adic heat equations, it is an open conjecture. By performing a Wick rotation on the heat equation, we obtain a Schrödinger equation, which is a continuous limit of a CTQMC. For general equations, this assertion is an open problem, but for the ones considered here, it was established in [26].
@@
We denote by $\mathbb{Z}_{p}$, where $p$ is a prime number, the unit ball of $\mathbb{Q}_{p}$, which can be visualize as an infinite rooted tree. By cutting this tree at level $l\geq 1$, we a finite rooted tree $G_{l}=\mathbb{Z}_{p}/p^{l}\mathbb{Z}_{p}$, with $l+1$ levels. The $p$-adic heat equations on the unit ball considered here have the form
@@
The above equation is a very particular example of a $p$-adic heat equation. In the general framework, the $p$-adic heat equations are deeply connected with models of complex hierarchical systems; they have been studied extensively in the last forty years; see, e.g. [13], [27]-[32], and the references therein.
@@
$\mathbb{Z}_{p}={\displaystyle\bigsqcup\limits_{I\in G_{l}}}\left(I+p^{l}\mathbb{Z}_{p}\right).$
We denote by $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ the characteristic function of the ball $I+p^{l}\mathbb{Z}_{p}$. We now construct a CTMC, with states $I\in G_{l}$, by defining the transition probability between states as
@@
Set $\boldsymbol{J}\varphi=J\ast\varphi-\varphi$, then the operators $\boldsymbol{J}$, $\mathrm{e}^{t\boldsymbol{J}}$, $\mathrm{e}^{it\boldsymbol{J}}$ are limits of finite dimensional operators $\boldsymbol{J}^{\left(l\right)}$, $\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}$, $\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}$, and consequently in the computation of $p_{r,v}(t)$, respectively of $\pi_{r,v}\left(t\right)$, we can use $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, respectively $Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$, and thus the construction of the above-mentioned Markov chains depends on a matrix $\boldsymbol{J}^{\left(l\right)}$. By suitably selecting this matrix, we construct the CTQWs based on adjacency matrices of graphs, see Appendix E.
As in [5], [3], the comparison between CTQMCs and CTMCs is based on the comparison between
@@
This work is organized into two parts. The first discusses the results and the simulations using a minimal mathematical background. The technical results were placed in appendices at the end of the paper.
@@ 2.1. The field of $p$-adic numbers
Remark 2.1.
The above formulation and results can be easily extended to the $N$-dimensional space $\mathbb{Q}_{p}^{N}$. We extend the $p$-adic norm to $\mathbb{Q}_{p}^{N}$ by taking
$x||_{p}:=\max_{1\leq i\leq N}|x_{i}|_{p},\qquad\text{for }x=(x_{1},\dots,x_{N})\in\mathbb{Q}_{p}^{N}.$
If we define $ord(x)=\min_{1\leq i\leq N}\{ord(x_{i})\}$, then $x||_{p}=p^{-ord(x)}$. The metric space $\left(\mathbb{Q}_{p}^{N},||\cdot||_{p}\right)$ is a complete ultrametric space. Ultrametric means that the norm $||\cdot||_{p}$ satisfies
[…]
A function $\varphi:\mathbb{Q}_{p}^{N}\rightarrow\mathbb{C}$ is called locally constant, if for any $a\in\mathbb{Q}_{p}^{N}$, there is an integer $l=l(a)$, such that $\varphi\left(a+x\right)=\varphi\left(a\right)$ for any $x||_{p}\leq p^{l}$. We denote by $\mathcal{D}(\mathbb{Q}_{p}^{N})$ the complex vector space of test functions. There is a natural integration theory so that $\int_{\mathbb{Q}_{p}^{N}}\varphi\left(x\right)d^{N}x$ gives a well-defined complex number. The measure $d^{N}x$ is the Haar measure of $\mathbb{Q}_{p}^{N}$.
@@ 2.2. Additive characters
Using that any non-zero $p$-adic number $x$ admits an expansion of the form $x=p^{ord(x)}\sum_{j=0}^{\infty}x_{j}p^{j}$, $x_{0}\neq 0$, the fractional part of $x\in\mathbb{Q}_{p}$, denoted $\{x\}_{p}$, is the rational number defined as
@@ 3.1. Initial value problems for $p$-adic Schrödinger equations on the unit ball
We set $\boldsymbol{H}:\mathcal{D}(\mathbb{Z}_{p})\rightarrow L^{2}(\mathbb{Z}_{p})$ to be a self-adjoint operator defined on the dense subspace $\mathcal{D}(\mathbb{Z}_{p})\hookrightarrow L^{2}(\mathbb{Z}_{p})$. Consider the following Cauchy problem
@@ 3.2. An orthonormal basis for $L^{2}(\mathbb{Z}_{p})$
and $\left\{\Psi_{rbk}\left(x\right)\right\}_{rbk}$ forms a complete orthonormal basis of $L^{2}(\mathbb{Q}_{p})$; see, e.g., [35, Theorems 9.4.5 and 8.9.3], or [13, Theorem 3.3].
@@ 3.3. Construction of CTQMCs I
Under the hypotheses 3.7 and 3.8, $\left[\pi_{s,r}\left(t\right)\right]_{s,r\in\mathbb{J}}$ is the transition matrix of a continuous-time quantum Markov chain with state space $\mathbb{J}$.
@@ 4.1. A class of $p$-adic heat equations
In Appendix B, using [40, Theorem 3.1], we show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)$ is the unique solution of (4.1), in a suitable function space, and that
@@ 7.1. Heat and Schrödinger equation attached to Bessel potentials
Then $\left\|J_{\alpha}\right\|_{1}=1$, $J_{\alpha}\left(|\xi|_{p}\right)\geq 0$, for $x\neq 0$, and $\widehat{J}_{\alpha}\left(|\xi|_{p}\right)=\left(\max\left\{1,\left|\xi\right|_{p}\right\}\right)^{-\alpha}$, see [41, Lemma 5.2].
@@ 8.2. Topology of $\mathbb{Q}_{p}$
As a topological space $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is totally disconnected, i.e., the only connected subsets of $\mathbb{Q}_{p}$ are the empty set and the points. A subset of $\mathbb{Q}_{p}$ is compact if and only if it is closed and bounded in $\mathbb{Q}_{p}$, see, e.g., [19, Section 1.3], or [35, Section 1.8]. The balls and spheres are compact subsets. Thus $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is a locally compact topological space.
@@ 8.7. Distributions
The $\mathbb{C}$-vector space $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ of all continuous linear functionals on $\mathcal{D}(\mathbb{Q}_{p})$ is called the Bruhat-Schwartz space of distributions. Every linear functional on $\mathcal{D}(\mathbb{Q}_{p})$ is continuous, i.e., $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ agrees with the algebraic dual of $\mathcal{D}(\mathbb{Q}_{p})$, see, e.g., [19, Chapter 1, VI.3, Lemma].

### v3

@@ 1. Introduction
In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement satisfies $\Delta x\geq L_{\text{Planck}}:=\sqrt{\frac{\hbar G}{c^{3}}}$, where $L_{\text{Planck}}\approx 10^{-33}$ cmcm is the Planck length. This implies that space-time is not an infinitely divisible continuum, [11]. Mathematically speaking, space-time must be a completely disconnected topological space. The ultrametric spaces are naturally completely disconnected. The field of $p$-adic numbers $\mathbb{Q}_{p}$ is a paramount example of an ultrametric space, with a very rich mathematical structure. There are several possible interpretations of the Bronstein inequality. One of them drives the loop quantum gravity. Volovich gave another interpretation of Bronstein’s inequality in the 80s. The inequality mentioned implies that real numbers cannot be used in models at the level of Planck’s length, because the Archimedean axiom, which appears naturally if we use real numbers, implies that lengths can be measured with arbitrary precision. Volovich proposed using $p$-adic numbers in physical models at the Planck scale, [12], see also [14]-[13, Chapter 6].
@@
In the Dirac-von Neumann formulation of quantum mechanics (QM), the states of a closed quantum system are vectors of an abstract complex Hilbert space $\mathcal{H}$, and the observables correspond to linear self-adjoint operators in $\mathcal{H}$. In $p$-adic QM, $\mathcal{H}=L^{2}(\mathbb{Q}_{p}^{N})$ or $\mathcal{H}=L^{2}(U)$, where $U\subset\mathbb{Q}_{p}^{N}$. This choice implies that the position vectors are $\mathbb{Q}_{p}^{N}$, while the time is a real variable. Using $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ as a model of space-time, implies assuming that the space ($\mathbb{Q}_{p}^{3}$) has a discrete nature: there are no continuous word lines joining two different points in the space. The assumption of the discreteness of space, which requires passing from $\mathbb{R}\times\mathbb{R}^{3}$ to $\mathbb{R}\times\mathbb{Q}_{p}^{3}$, is incompatible with special and general relativity. We warn the reader that there are several different types of $p$-adic QM; for instance, if the time is assumed to be a $p$-adic variable, the QM obtained radically differs from the one considered here. To the best of our knowledge, $p$-adic QM started in the 1980s under the influence of Vladimirov and Volovich, [15]. The literature about $p$-adic QM is pervasive, and here we cited just a few works, [15]-[25].
@@
In [23], the first author showed that a large class of $p$-adic Schrödinger equations are the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices were introduced. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. The final goal of the mentioned work was to contribute in the understanding of the role of the hypothesis of the discreteness of space in the foundations of quantum mechanics. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. The paradigm asserting that the universe is not locally real implies that $p$-adic QM allows realism.
@@
In this work, we study a large class of $p$-adic Schrödinger equations and the associated CTQMCs by comparing them with $p$-adic heat equations and the associated CTMCs. This work is motivated by [5], where a particular class CTQWs was compared with their classical counterparts. Here, the comparison is done by a mathematical study of the mentioned equations, which requires, for instance, solving the initial value problems attached to the mentioned equations, and through numerical simulations. A $p$-adic heat equation is a continuous (or scaling) limit of a CTMC. This assertion has been verified for some classes of equations, see [23], which include the one considered here. But, for general $p$-adic heat equations, it is an open conjecture. By performing a Wick rotation on a heat equation, we obtain a Schrödinger equation, which is a continuous limit of a CTQMC. For general equations, this assertion is an open problem, but for the ones considered here, it was established in [23].
@@
We denote by $\mathbb{Z}_{p}$, where $p$ is a prime number, the unit ball of $\mathbb{Q}_{p}$, which can be visualize as an infinite rooted tree. The $p$-adic heat equations on the unit ball considered here have the form
@@
The above equation is a particular example of a $p$-adic heat equation. In the general framework, the $p$-adic heat equations are deeply connected with models of complex hierarchical systems; they have been studied extensively in the last forty years; see, e.g. [26]-[32], and the references therein.
@@
$\mathbb{Z}_{p}={\displaystyle\bigsqcup\limits_{I\in G_{l}}}\left(I+p^{l}\mathbb{Z}_{p}\right),$
where $G_{l}=\mathbb{Z}_{p}/p^{l}\mathbb{Z}_{p}$ is a finite rooted tree with $l+1$ levels. We denote by $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ the characteristic function of the ball $I+p^{l}\mathbb{Z}_{p}$. We now construct a CTMC, with states $I\in G_{l}$, by defining the transition probability between states as
@@
Set $\boldsymbol{J}\varphi=J\ast\varphi-\varphi$, then the operators $\boldsymbol{J}$, $\mathrm{e}^{t\boldsymbol{J}}$, $\mathrm{e}^{it\boldsymbol{J}}$ are limits of finite dimensional operators $\boldsymbol{J}^{\left(l\right)}$, $\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}$, $\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}$, and consequently in the computation of $p_{r,v}(t)$, respectively of $\pi_{r,v}\left(t\right)$, we can use $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, respectively $Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$, and thus the construction of the above-mentioned Markov chains depends on a matrix $\boldsymbol{J}^{\left(l\right)}$. By suitably selecting this matrix, we construct the CTQWs based on graph adjacency matrices; see Appendix E.
Following [5], [3], the comparison between CTQMCs and CTMCs is based on the comparison between
@@
2. The road map
The paper was written for a large audience, including physicists, mathematicians, and computer scientists. For this reason, the paper was written so that most of the technical (mathematical) results can be avoided on a first reading. This work is organized into two parts. The first discusses the results and simulations; only some mathematical results are discussed due to their relevance in the construction of CTQWs. But most of the technical results were placed in appendices at the end of the paper. The appendices contain a detailed (as much as possible) presentation of the mathematical results, intended for mathematically inclined readers.
This paper is part of a research effort led by the first author to understand the connection between the foundations of QM from the perspective of the hypothesis that the physical space has a discrete nature at short distances, [21]-[25]. Remarkably, this investigation is connected with quantum computing.
In [23], the first author established that the $p$-adic QM is a non-local theory that admits realism, by showing that a large class of $p$-adic Schrödinger equations are continuous versions of CTQWs. A very detailed discussion of the foundations of QM and its connections with $p$-adic QM is presented in [23], and [25]. Here, we just mention that the connection between the hypothesis that the physical space has a discrete nature and CTQWs was already discussed in [23, p. 2]: “Since $p$-adic QM is nonlocal, this theory admits a reality independent of the observer: quantum particles exist, and they move randomly in a configuration space of type $\mathbb{R}\times\mathcal{X}^{N}$, this means that the quantum motion can be modeled as a Markov process in $\mathbb{R}\times\mathcal{X}^{N}$. Intuitively, the CTQWs introduced here are approximations of the quantum motion mentioned.” Here, $\mathcal{X}$ is a mathematical model of a discrete space.
After [23], a natural question is whether a new theory for CTQWS based on $p$-adic analysis is possible. This paper constitutes the first step in this direction. Here, we present a new theory for CTQWs, including simulations and comparisons. The results are presented in a mathematically rigorous manner; we aim to provide a short, almost self-contained paper. We warn the reader that this taks is difficult, the theory of $p$-adic heat equations is already very extensive and technical; see, for instance, [28]. The heat equations used in this paper were studied in [29] and [30].
The first author has conjectured that every heat equation is a continuous limit of a continuous-time random walk, and that the corresponding Schrödinger equation is a continuous limit of a CTQW. The study of each CTQW goes through the study a particular $p$-adic Schrödinger equation and its discretization. Here we consider only a partcicular class of heat and Schrödinger equations whose discretizations are easily computed.
There are several different constructions of CTQMCs based on $p$-adic Schrödinger equations. In Sections 4.3, 4.4, we review these constructions and introduce new ones. In Section 7, we present several numerical simulations comparing classical random walks against their quantum counterparts. This requires selecting a kernel $J$; we use the Bessel potentials $J_{\alpha}\left(\left|x\right|_{p}\right)$ because they have been extensively studied, and many useful properties and formulas are known. The simulations show the behavior of the transition probability $p_{I,J}(t)$ between state $J$ and $I$ at time $t$, and the corresponding quantum version $\pi_{I,J}(t)$. The simulations follow the approach described in [3], [5], in which the authors studied a specific network and compared $p_{I}^{\text{sta}}$ and $\chi_{I,J}$. The inequality $p_{I}^{\text{sta}}$ << $\chi_{I,J}$ is interpreted as the fact that quantum random walks have greater computational power than classical ones. The rationale behind this assertion is that the entries in the matrix $\chi_{I,J}$ are spread out around the diagonal, showing that the walker can perform random walks around each state, while $p_{I}^{\text{sta}}$ is a diagonal matrix, which means that the probability that the walker jumps between two different states is zero. This phenomenon also occurs in our simulations. The last section (Discussion) provides a summary of the results and the contributions of the present paper in our research program on the foundations of QM, space discreteness, and CTQWs.
@@ 3.2. Additive characters
Using that any non-zero $p$-adic number $x$ admits an expansion of the form $x=p^{ord(x)}\sum_{j=0}^{\infty}x_{j}p^{j}$, $x_{0}\neq 0$, the fractional part of $x\in\mathbb{Q}_{p}$, denoted as $\{x\}_{p}$, is the rational number defined as
@@ 4.1. Initial value problems for $p$-adic Schrödinger equations on the unit ball
We set $\boldsymbol{H}:\mathcal{D}(\mathbb{Z}_{p})\rightarrow L^{2}(\mathbb{Z}_{p})$ to be a self-adjoint operator defined on the dense subspace $\mathcal{D}(\mathbb{Z}_{p})\hookrightarrow L^{2}(\mathbb{Z}_{p})$. Consider the Cauchy problem
@@ 4.2. An orthonormal basis for $L^{2}(\mathbb{Z}_{p})$
and $\left\{\Psi_{rbk}\left(x\right)\right\}_{rbk}$ forms a complete orthonormal basis of $L^{2}(\mathbb{Q}_{p})$; see, e.g., [35, Theorems 9.4.5 and 8.9.3], or [26, Theorem 3.3].
@@ 4.3. Construction of CTQMCs I
Under the hypotheses 4.7 and 4.8, $\left[\pi_{s,r}\left(t\right)\right]_{s,r\in\mathbb{J}}$ is the transition matrix of a continuous-time quantum Markov chain with state space $\mathbb{J}$.
@@ 5.1. A class of $p$-adic heat equations
In Appendix B, using [29, Theorem 3.1], we show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)$ is the unique solution of (5.1), in a suitable function space, and that
@@ 8.1. Heat and Schrödinger equation attached to Bessel potentials
Then $\left\|J_{\alpha}\right\|_{1}=1$, $J_{\alpha}\left(|\xi|_{p}\right)\geq 0$, for $x\neq 0$, and $\widehat{J}_{\alpha}\left(|\xi|_{p}\right)=\left(\max\left\{1,\left|\xi\right|_{p}\right\}\right)^{-\alpha}$, see [39, Lemma 5.2].
@@ 9.2. Topology of $\mathbb{Q}_{p}$
As a topological space $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is totally disconnected, i.e., the only connected subsets of $\mathbb{Q}_{p}$ are the empty set and the points. A subset of $\mathbb{Q}_{p}$ is compact if and only if it is closed and bounded in $\mathbb{Q}_{p}$, see, e.g., [16, Section 1.3], or [35, Section 1.8]. The balls and spheres are compact subsets. Thus $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is a locally compact topological space.
@@ 9.7. Distributions
The $\mathbb{C}$-vector space $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ of all continuous linear functionals on $\mathcal{D}(\mathbb{Q}_{p})$ is called the Bruhat-Schwartz space of distributions. Every linear functional on $\mathcal{D}(\mathbb{Q}_{p})$ is continuous, i.e., $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ agrees with the algebraic dual of $\mathcal{D}(\mathbb{Q}_{p})$, see, e.g., [16, Chapter 1, VI.3, Lemma].